A railroad car of 2000 kg coasting at 3.0 m/s overtakes and locks together with an identical car coasting on the same track in the same direction at 2.0 m/s. What is the speed of the cars after the lock together?
Answer #1

m = mass of each railroad car = 2000 kg

v1 = initial velocity of faster car before the collision = 3 m/s

v2 = initial velocity of slower car before the collision = 2 m/s

v = final velocity of the combination of two car after collision = ?

since the two cars lock together after collision, hence the collision is inelastic.

Using conservation of momentum , we get

m v1 + m v2 = m v + m v

v1 + v2 = 2 v

inserting the values

3 + 2 = 2 v

v = 5/2

v = 2.5 m/s